# Permutation-Invariant and Orientation-Aware Dataset Distillation for 3D Point Clouds

### v1

@@ 5 Conclusion
We proposed a permutation-invariant and orientation-aware dataset distillation framework for 3D point clouds. To alleviate the inconsistency of point ordering between compared 3D models, we devise a Permutation Invariant Distribution Matching (PIDM) loss where the sorted features are compared. We also employed learnable parameters of rotation angles to estimate optimal orientation of synthetic models. The geometric structures and orientation of the synthetic models are jointly optimized during the dataset distillation. Experimental results evaluated on four widely used benchmark datasets of ModelNet10 [24], ModelNet40 [24], ShapeNet [4], and ScanObjectNN [19] demonstrate that the proposed method outperforms existing distillation methods while maintaining strong cross-architecture generalization. Furthermore, we also validated the effectiveness of each component of the proposed method through extensive ablation studies.

### v2

@@ 5 Conclusion
We proposed a semantically-aligned and orientation-aware dataset distillation framework for 3D point clouds. To alleviate the inconsistency of point ordering between compared 3D objects, we devised a Semantically Aligned Distribution Matching loss where the sorted features in each channel are compared. We also employed learnable parameters of the rotation angles to estimate the optimal orientation of synthetic objects. The geometric structures and orientations of the synthetic objects are jointly optimized during the dataset distillation. Experimental results evaluated on the four widely used benchmark datasets of ModelNet10 [24], ModelNet40 [24], ShapeNet [4], and ScanObjectNN [19] demonstrate that the proposed method outperforms existing distillation methods while maintaining strong cross-architecture generalization. Furthermore, we also validated the effectiveness of each component of the proposed method through extensive ablation studies.